# Entanglement harvesting and curvature of entanglement: A modular operator approach

### v1

@@ Abstract
An operator-algebraic framework based on Tomita-Takesaki modular theory is used to study aspects of quantum entanglement via the application of the modular conjugation operator $J$. The entanglement structure of quantum fields is studied through the protocol of entanglement harvesting whereby by quantum correlations evolve through the time evolution of qubit detectors coupled to a Bosonic field. Modular conjugation operators are constructed for Unruh-Dewitt type qubits interacting with a scalar field such that initially unentangled qubits become entangled. The entanglement harvested in this process is directly quantified by an expectation value involving $J$ offering a physical application of this operator. The modular operator formalism is then extended to the Markovian open system dynamics of coupled qubits by expressing entanglement monotones as functionals of a state $\rho$ and its modular reflection $J\rho J$. The second derivative of such functionals with respect to an external coupling parameter, termed the curvature of entanglement, provides a natural measure of entanglement sensitivity. At points of modular self-duality, the curvature of entanglement coincides with the quantum Fisher information measure. These results demonstrate that the modular conjugation operator $J$ captures both the harvesting of entanglement from quantum fields and the curvature of entanglement in coupled qubit dynamics providing parallel modular structures that connect these systems.
@@ I Introduction
The remainder of this paper is organized as follows. In Sec. II, we review the some elements of Tomita–Takesaki modular theory and define the modular conjugation operator and modular reflection that underpin our framework. In Sec. III, we apply this structure to entanglement harvesting using Unruh–DeWitt detectors interacting with a scalar quantum field, and show how a modular expectation value quantifies the harvested entanglement. In Sec. IV, we reformulate Wootters’ concurrence for mixed states as a modular functional $f(\rho,J\rho J)$ in order to reinterpret entanglement through modular reflection. In Sec. V, we introduce the concept of modular curvature of entanglement as the second derivative of such functionals with respect to a coupling parameter. In Sec. V-A, we revisit the coupled qubit model and recover the curvature of entanglement behavior previously studied in [3], now interpreted through our operator-algebraic framework. We show that this curvature coincides with quantum Fisher information at points of modular self-duality. In Sec. V-B, we discuss the second-order expansion of relative entropy, and the emergence of quantum Fisher information as a curvature measure tied to modular self-duality. Finally, in Sec. VI, we outline future directions, including a broader class of modular functionals and conjectured monotone Riemannian metrics defined via the modular conjugation operator, thereby embedding entanglement curvature into a unified operator-algebraic framework. Additional technical derivations are included in the Appendix.
@@ II Tomita–Takesaki Modular Operators
where G(xy)G(x-y) is a propagator such as a Pauli-Jordon Green’s function. This condition ensures that the smeared field operators commute,
@@
The term $2|X|$, arising entirely from field-induced off-diagonal coherence, is precisely the modular expectation value $\braket{\Psi_{f}|J_{AB}\otimes I_{F}|\Psi_{f}}$. This quantity captures the entanglement generated between the detectors due to spacelike correlations in the quantum field, as encoded by the Tomita–Takesaki modular conjugation operator.
@@ IV Modular Formulation of Concurrence for Mixed States
To extend Tomita–Takesaki theory to mixed states, we employ the GNS construction associated with the state $\rho$. Let $\rho$ be a density operator acting on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, and let $J$ denote the Tomita–Takesaki modular conjugation operator with respect to a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ and a cyclic separating vector $|\Omega\rangle$ representing $\rho$. If $\rho$ is a pure state and is maximally entangled, its concurrence satisfies $C(\rho)=1$. More generally, if $\rho=J\rho J$ but $\rho$ is mixed, then the concurrence is bounded above by unity but not necessarily at maximal entanglement. This establishes that exact modular self-duality is a necessary condition for maximal entanglement in pure states and positions concurrence as a functional that quantifies its violation. The physical content of this construction is reinforced by its connection to Bell inequalities. In particular, the algebraic concurrence defined above satisfies the maximal Bell–CHSH inequality violation:
@@ V Curvature of Entanglement
where the inner products can be explicitly written as traces. Since the modular conjugation operator $J$ is antiunitary and independent of the coupling parameter $g$, we have
$\frac{\partial(J\rho J)}{\partial g}=J\left(\frac{\partial\rho}{\partial g}\right)J,\quad\frac{\partial^{2}(J\rho J)}{\partial g^{2}}=J\left(\frac{\partial^{2}\rho}{\partial g^{2}}\right)J.$ (34)
@@ V.2 Relative Entropy, Curvature, and Modular Sensitivity
The relative entropy between two quantum states $\rho$ and $\sigma$ is defined as
@@
and serves as a fundamental measure of distinguishability in quantum information theory. It plays a key role in thermodynamics, quantum hypothesis testing, and information-theoretic bounds on state conversion [23, 24, 1].
In the operator-algebraic formulation of modular theory, this quantity acquires a deeper structural interpretation. Specifically, using the Tomita–Takesaki relative modular operator $\Delta_{\sigma,\rho}$, one obtains the identity
@@
where $|\Omega_{\rho}\rangle$ is the cyclic and separating GNS vector associated with $\rho$. This formulation, due to Araki [25], generalizes relative entropy to arbitrary von Neumann algebras and connects it directly to the modular flow generated by the pair $(\rho,\sigma)$. The operator $\log\Delta_{\sigma,\rho}$ functions as a relative modular Hamiltonian, and its expectation value measures the entropic asymmetry between $\rho$ and $\sigma$ under Tomita–Takesaki evolution [28, 1].
Let $\rho(g)$ be a smoothly varying family of density operators parametrized by a real coupling parameter $g$. Consider a perturbed version of $\rho(g)$ under a small completely positive trace-preserving (CPTP) map $\mathcal{E}_{\varepsilon}$, such that $\mathcal{E}_{0}=\mathrm{id}$. Then the relative entropy between the unperturbed and perturbed states satisfies the expansion
@@
where $F(g)$ denotes the quantum Fisher information associated with the family $\rho(g)$. This second-order behavior reveals that $F(g)$ quantifies the local curvature of the quantum state manifold with respect to the relative entropy. This result has been rigorously formalized in terms of channel reversibility by Fawzi and Renner in [26]. They established that small relative entropy implies the existence of a nearly optimal recovery map, specifically the Petz recovery map, while Sutter, Tomamichel, and Berta [27] demonstrated tight second-order expansions that quantify the extent of irreversibility under noise. To leading order, a small value of $F(g)$ indicates that the relative entropy between $\rho(g)$ and $\rho(g+\varepsilon)$ grows slowly with $\varepsilon$, implying that the state departs only gradually from being reversible under the relevant CPTP channel, and hence remains close to being recoverable by the Petz map. When $F(g)$ is large, the breakdown of reversibility is sharply reflected in the curvature of the information geometry: the relative entropy between $\rho(g)$ and $\rho(g+\varepsilon)$ grows rapidly with $\varepsilon$, indicating strong sensitivity to perturbations and a correspondingly diminished effectiveness of the Petz recovery map.
We emphasize that this geometric curvature has an operator-algebraic analog in the entanglement curvature $\mathcal{K}_{E}(g)$ revealing that the modular conjugation operator $J$ encodes the same local sensitivity as the quantum Fisher information but framed entirely within the algebraic structure of Tomita–Takesaki theory. This is completely parallel with the construction of relative entropy being constructed with the relative modular operator $\Delta_{\sigma,\rho}$.
While quantum Fisher information arises from local statistical distinguishability and optimal parameter estimation, modular frame curvature reflects the sensitivity of entanglement structure under modular conjugation. Their coincidence at self-duality points gives precise geometric and operational meaning to the modular symmetry condition $\rho=J\rho J$. At such points, the entanglement curvature $\mathcal{K}_{E}(g)$ coincides with the negative quantum Fisher information, i.e., $\mathcal{K}_{E}(g)=-F(g)$, providing a direct modular signature of reversibility.
Away from self-duality, however, these quantities diverge. The curvature $\mathcal{K}_{E}(g)$ no longer tracks the Fisher information but instead quantifies the breakdown of modular invariance under parameter evolution. Specifically, it measures the second-order deviation from the modular fixed-point condition $\rho(g)=J\rho(g)J$, and hence the rate at which entanglement becomes asymmetric under conjugation. In this regime, $\mathcal{K}_{E}(g)$ continues to capture meaningful geometric information—one that persists even when quantum Fisher information becomes difficult to interpret due to singularities, rank-deficiency, or non-Markovianity.
@@ VI Conclusion and Future Research
Taken together, these findings suggest that modular operator theory offers a promising framework for relating entanglement structure, curvature, and recoverability within quantum information theory. We expect that the operator-algebraic techniques developed here extend beyond the specific models studied, providing a modular toolkit for analyzing entanglement dynamics, quantum channel reversibility, and information geometry. Potential applications include quantum metrology, quantum resource theories, and the structure of entanglement in quantum field theory and holography.

### v2

@@ Abstract
An operator-algebraic framework based on Tomita-Takesaki modular theory is used to study aspects of quantum entanglement via the application of the modular conjugation operator $J$. The entanglement structure of quantum fields is studied through the protocol of entanglement harvesting whereby quantum correlations evolve through the time evolution of qubit detectors coupled to a Bosonic field. Modular conjugation operators are constructed for Unruh-Dewitt type qubits interacting with a scalar field such that initially unentangled qubits become entangled. The entanglement harvested in this process is directly quantified by an expectation value involving $J$ offering a physical application of this operator. The modular operator formalism is then extended to the Markovian open system dynamics of coupled qubits by expressing entanglement monotones as functionals of a state $\rho$ and its modular reflection $J\rho J$. The second derivative of such functionals with respect to an external coupling parameter, termed the curvature of entanglement, provides a natural measure of entanglement sensitivity. At points of modular self-duality, the curvature of entanglement coincides with the quantum Fisher information measure. These results demonstrate that the modular conjugation operator $J$ captures both the harvesting of entanglement from quantum fields and the curvature of entanglement in coupled qubit dynamics providing parallel modular structures that connect these systems.
@@ I Introduction
In what follows, we work within the setting of type I von Neumann algebras corresponding to finite-dimensional Hilbert spaces of the effective detector models used here. In this context, density operators and tensor-product factorizations $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ are well defined ensuring that modular conjugation can be represented by a fixed antiunitary operator acting on a separable Hilbert space and permitting explicit constructions of functionals of the form $f(\rho,J\rho J)$.
@@
The remainder of this paper is organized as follows. In Sec. II, we review some elements of Tomita–Takesaki modular theory and define the modular conjugation operator and modular reflection that underpin our framework. In Sec. III, we apply this structure to entanglement harvesting using Unruh–DeWitt detectors interacting with a scalar quantum field, and show how a modular expectation value quantifies the harvested entanglement. In Sec. IV, we reformulate Wootters’ concurrence for mixed states as a modular functional $f(\rho,J\rho J)$ in order to reinterpret entanglement through modular reflection. In Sec. V, we introduce the concept of modular curvature of entanglement as the second derivative of such functionals with respect to a coupling parameter. In Sec. V-A, we revisit the coupled qubit model and recover the curvature of entanglement behavior previously studied in [3], now interpreted through our operator-algebraic framework. We show that this curvature coincides with quantum Fisher information at points of modular self-duality. In Sec. V-B, we discuss the second-order expansion of relative entropy, and the emergence of quantum Fisher information as a curvature measure tied to modular self-duality. Finally, in Sec. VI, we outline future directions, including a broader class of modular functionals and conjectured monotone Riemannian metrics defined via the modular conjugation operator, thereby embedding entanglement curvature into a unified operator-algebraic framework. Additional technical derivations are included in the Appendix.
@@ II Tomita–Takesaki Modular Operators
where G(xy)G(x-y) is a propagator such as a Pauli-Jordan Green’s function. This condition ensures that the smeared field operators commute,
@@
The term $2|X|$, arising entirely from field-induced off-diagonal coherence, is precisely the modular expectation value $\braket{\Psi_{f}|J_{AB}\otimes I_{F}|\Psi_{f}}$. This quantity captures the entanglement generated between the detectors due to spacelike correlations in the quantum field, as encoded by the Tomita–Takesaki modular conjugation operator. Although $J_{AB}$ is antiunitary and therefore not a Hermitian observable, the overlap $\braket{\Psi_{f}|J_{AB}\otimes I_{F}|\Psi_{f}}$ can be interpreted operationally as an interference amplitude between the final state and its modularly reflected counterpart. Its modulus, $2|X|$, quantifies the visibility of coherence between excitation pathways that are exchanged under detector interchange and complex conjugation. Interferometric and weak-measurement protocols have been shown to directly access such overlap amplitudes and coherences, without requiring full state tomography [17, 18, 19, 20]. In this sense, the quantity $2|X|$ plays the same operational role as measurable coherence and fidelity estimators in resource-theoretic formulations of quantum coherence [21]. Thus, while $J_{AB}$ itself is not directly measurable, the interference pattern it defines possesses clear operational meaning and coincides with the concurrence’s coherence term. These same interference visibilities appear in recent optical entanglement-harvesting experiments and photonic state-overlap measurements suggesting possible realizations of the modular overlap in cavity-QED or circuit-QED platforms.
@@ IV Modular Formulation of Concurrence for Mixed States
To extend Tomita–Takesaki theory to mixed states, we employ the GNS construction associated with the state $\rho$. Let $\rho$ be a density operator acting on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, and let $J$ denote the Tomita–Takesaki modular conjugation operator with respect to a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ and a cyclic separating vector $|\Omega\rangle$ representing $\rho$. If $\rho$ is a pure state and is maximally entangled, its concurrence satisfies $C(\rho)=1$. More generally, if $\rho=J\rho J$ but $\rho$ is mixed, then the concurrence is bounded above by unity but not necessarily at maximal entanglement. This establishes that exact modular self-duality is a necessary condition for maximal entanglement in pure states and positions concurrence as a functional that quantifies its violation.
It is important to distinguish between the pure- and mixed-state cases when discussing modular self-duality. For pure states, modular self-duality $\rho=J\rho J$ is both a necessary and sufficient condition for maximal entanglement, since it coincides with equal Schmidt coefficients and $C(\rho)=1$.
However, for mixed states, modular self-duality remains necessary but not sufficient for maximal entanglement. In particular, any state that is diagonal in a basis symmetric under $J$ (such as the maximally mixed state) trivially satisfies this equality while possessing zero concurrence. Maximal entanglement arises only when self-duality holds together with additional spectral and support conditions—namely, when the support of $\rho$ lies within the Bell subspace generated by the eigenvectors of $R=\rho J\rho J$ with non-zero off-diagonal coherence. Equivalently, for rank-one (pure) or rank-two states with equal Schmidt coefficients, modular self-duality coincides with $C(\rho)=1$. This distinction clarifies that $J\rho J=\rho$ characterizes the fixed-point symmetry, while concurrence quantifies the degree of its violation.
The physical content of this construction is reinforced by its connection to Bell inequalities. In particular, the algebraic concurrence defined above for two qubit states satisfies the maximal Bell-CHSH inequality violation,
@@
In the finite-dimensional setting considered here, the modular conjugation operator is defined within a fixed representation of the type I von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$. The operator $J_{\mathrm{QIT}}=(\sigma_{y}\otimes\sigma_{y})K$ acts as an antiunitary reflection that exchanges the two subsystems and performs complex conjugation in the computational basis. As detailed in Appendix B, any other representation obtained by local unitaries $U_{A}\otimes U_{B}$ yields an equivalent conjugation $J^{\prime}=(U_{A}\otimes U_{B})J_{\mathrm{QIT}}(U_{A}^{\dagger}\otimes U_{B}^{\dagger})$, so the modular structure is representation dependent but algebraically fixed. Once a computational representation is chosen, $J$ therefore remains invariant under all physical operations within that representation. This representational invariance underlies the subsequent analysis, where the same modular conjugation is used to define curvature functionals of the density operator.
@@ V Curvature of Entanglement
where the inner products can be explicitly written as traces.
As established in Sec. IV, $J$ acts within a fixed computational representation of the type I algebra and is therefore treated as constant when evaluating parameter derivatives. Since the modular conjugation operator $J$ is antiunitary and representation-dependent but not dynamically dependent on the coupling parameter $g$, we have
$\frac{\partial(J\rho J)}{\partial g}=J\!\left(\frac{\partial\rho}{\partial g}\right)\!J,\qquad\frac{\partial^{2}(J\rho J)}{\partial g^{2}}=J\!\left(\frac{\partial^{2}\rho}{\partial g^{2}}\right)\!J.$ (34)
@@ V.2 Relative Entropy, Curvature, and Modular Sensitivity
In quantum information theory, the relative entropy between two faithful normal states $\rho$ and $\sigma$ is defined as
@@
and serves as a fundamental measure of statistical distinguishability between quantum states. It plays a central role in quantum thermodynamics, hypothesis testing, and information-theoretic bounds on state conversion [28, 29, 1]. Equation (60) is well defined whenever $\rho$ and $\sigma$ are density operators on a finite-dimensional Hilbert space or, more generally, normal states on a type I von Neumann algebra.
For type III von Neumann algebras—ubiquitous in algebraic quantum field theory—no normal trace exists, and hence the expression (60) cannot be defined directly. In this setting, the operator-algebraic framework of Tomita–Takesaki modular theory provides a general, trace-independent formulation of relative entropy. Let $\mathfrak{M}\subset\mathcal{B}(\mathcal{H})$ denote a von Neumann algebra of bounded operators acting on a Hilbert space $\mathcal{H}$, closed in the weak operator topology and containing the identity operator. The algebra $\mathfrak{M}$ represents the physical observables associated with a given system or spacetime region. A faithful normal state $\rho$ on $\mathfrak{M}$ is represented by a cyclic and separating vector $|\Omega_{\rho}\rangle\in\mathcal{H}$ satisfying
$\rho(A)=\langle\Omega_{\rho}|A|\Omega_{\rho}\rangle,\qquad A\in\mathfrak{M}.$ (61)
The pair $(\mathfrak{M},|\Omega_{\rho}\rangle)$ thus forms the standard representation of the algebra. Following Araki [30], one associates to two such faithful normal states $\rho$ and $\sigma$ the relative modular operator $\Delta_{\sigma,\rho}$, defined by
$\Delta_{\sigma,\rho}=S_{\sigma,\rho}^{\dagger}S_{\sigma,\rho},$ (62)
where the relative Tomita operator $S_{\sigma,\rho}$ acts on the dense domain
$\mathcal{D}(S_{\sigma,\rho})=\{A|\Omega_{\rho}\rangle:A\in\mathfrak{M}\},$ (63)
according to
$S_{\sigma,\rho}\,A|\Omega_{\rho}\rangle=A^{\dagger}|\Omega_{\sigma}\rangle.$ (64)
Here, $|\Omega_{\sigma}\rangle$ is the GNS vector corresponding to the state $\sigma$, also cyclic and separating for $\mathfrak{M}$.
The relative modular operator is characterized by the quadratic form
$\langle\psi|\Delta_{\sigma,\rho}\phi\rangle=\langle S_{\sigma,\rho}\psi|S_{\sigma,\rho}\phi\rangle,\qquad\forall\,\psi,\phi\in\mathcal{D}(S_{\sigma,\rho}).$ (65)
This construction requires neither a trace nor a matrix representation, making it valid for general type III factors. With this operator, Araki’s relative entropy is expressed as
@@
which generalizes the finite-dimensional expression to arbitrary von Neumann algebras. In this formulation, the operator $\log\Delta_{\sigma,\rho}$ plays the role of the relative modular Hamiltonian in the Tomita–Takesaki framework [30, 33, 1]. This expression remains meaningful in type III algebras, where $\Delta_{\sigma,\rho}$ has a continuous spectrum and the trace-based definition is not applicable.
In the special case of type I von Neumann algebras, the construction reduces to the familiar finite-dimensional form. One obtains the concrete realization
$\Delta_{\sigma,\rho}=\sigma\otimes\rho^{-1},$ (67)
acting on the Hilbert–Schmidt space with GNS vector $|\Omega_{\rho}\rangle=\sum_{i}\sqrt{\lambda_{i}}\,|\phi_{i}\rangle\otimes|\phi_{i}\rangle$, where $\rho=\sum_{i}\lambda_{i}|\phi_{i}\rangle\langle\phi_{i}|$. In this case, Eq. (66) reproduces the standard relative entropy found in quantum information theory.
The explicit constructions of the Tomita operator $S_{\rho}$, modular conjugation $J_{\rho}$, and modular operator $\Delta_{\rho}$, together with their modular-flow relations, are summarized in Appendix A. A systematic definition of “modular concurrence” or “entanglement curvature” in genuine type III settings would require replacing density-operator expressions by modular pairs of states such as $(\rho,J\rho J)$ interpreted through relative modular operators or via the split property, which provides an effective type I inclusion between nested local algebras. While a full generalization along these lines lies beyond the present scope, the modular-theoretic relations shown here form a consistent operator-algebraic foundation from which such type III extensions can be constructed.
Let $\rho(g)$ denote a smoothly varying family of density operators parametrized by a real coupling parameter $g$. Consider a perturbed version of $\rho(g)$ under a small completely positive trace-preserving (CPTP) map $\mathcal{E}_{\varepsilon}$ such that $\mathcal{E}_{0}=\mathrm{id}$. The relative entropy between the unperturbed and perturbed states then satisfies the second-order expansion
@@
where $F(g)$ denotes the quantum Fisher information associated with the family $\rho(g)$. This second-order behavior reveals that $F(g)$ quantifies the local curvature of the quantum state manifold with respect to the relative entropy. Fawzi and Renner [31] rigorously established that small relative entropy implies the existence of an almost optimal recovery channel—specifically the Petz map—while Sutter, Tomamichel, and Berta [32] obtained tight second-order expansions quantifying the degree of irreversibility under noise. To leading order, a small value of $F(g)$ indicates that the relative entropy between $\rho(g)$ and $\rho(g+\varepsilon)$ grows slowly with $\varepsilon$, signifying that the state remains nearly reversible under the relevant CPTP channel and is well approximated by the Petz recovery map. Conversely, a large $F(g)$ signals that reversibility breaks down sharply—the relative entropy grows rapidly with $\varepsilon$, and the state exhibits strong sensitivity to perturbations.
This geometric curvature has an operator-algebraic counterpart in the entanglement curvature $\mathcal{K}_{E}(g)$, where the modular conjugation operator $J$ encodes the same local sensitivity captured by the quantum Fisher information, but expressed entirely within the Tomita–Takesaki algebraic framework. Just as relative entropy is defined through the relative modular operator $\Delta_{\sigma,\rho}$, the curvature $\mathcal{K}_{E}(g)$ reflects second-order variations of modular structure and hence the “modular sensitivity” of the entanglement geometry.
While the quantum Fisher information quantifies local statistical distinguishability and optimal parameter estimation, the modular curvature measures the responsiveness of the entanglement structure under modular conjugation. At self-duality points satisfying $\rho=J\rho J$, these two notions coincide, yielding $\mathcal{K}_{E}(g)=-F(g)$; the equality provides a precise modular signature of reversibility. Away from self-duality, however, the two curvatures diverge: $\mathcal{K}_{E}(g)$ no longer tracks $F(g)$ but instead measures the degree of modular asymmetry and the rate at which the state departs from the fixed-point condition $\rho(g)=J\rho(g)J$. Even in regimes where $F(g)$ loses its operational meaning—due to singularities, rank deficiency, or non-Markovian dynamics—the modular curvature $\mathcal{K}_{E}(g)$ continues to furnish a robust geometric indicator of entanglement sensitivity.
@@ VI Conclusion and Future Research
Taken together, these findings suggest that modular operator theory offers a promising framework for relating entanglement structure, curvature, and recoverability within quantum information theory. We expect that the operator-algebraic techniques introduced here extend beyond the specific models studied, providing a modular toolkit for analyzing entanglement dynamics, quantum channel reversibility, and information geometry. Potential applications include quantum metrology [40], quantum resource theories [41], and the structure of entanglement in quantum field theory [37] and holography [42]. Moreover, as noted in Sec. V(B), the connection to Araki’s relative modular operator $\Delta_{\sigma,\rho}$ and relative entropy provides a pathway toward generalizing the present framework to type III von Neumann algebras, where density-matrix representations are unavailable. Within algebraic quantum field theory, such generalizations would naturally invoke the split property to realize approximate type I factorizations and to characterize modular sensitivity in local algebras. These extensions connecting modular concurrence and entanglement curvature to the operator-algebraic foundations of field theory will be explored in future work.